### v1

# Physics-Informed Singular-Value Learning for Cross-Covariances Forecasting in Financial Markets

Abstract
A new wave of work on covariance cleaning and nonlinear shrinkage has delivered asymptotically optimal analytical solutions for large covariance matrices. Building on this progress, these ideas have been generalized to empirical cross–covariance matrices, whose singular-value shrinkage characterizes comovements between one set of assets and another. Existing analytical cross-covariance cleaners are derived under strong stationarity and large-sample assumptions, and they typically rely on mesoscopic regularity conditions such as bounded spectra; macroscopic common modes (e.g., a global market factor) violate these conditions. When applied to real equity returns, where dependence structures drift over time and global modes are prominent, we find that these theoretically optimal formulas do not translate into robust out-of-sample performance. We address this gap by designing a random-matrix-inspired neural architecture that operates in the empirical singular-vector basis and learns a nonlinear mapping from empirical singular values to their corresponding cleaned values. By construction, the network can recover the analytical solution as a special case, yet it remains flexible enough to adapt to non-stationary dynamics and mode-driven distortions. Trained on a long history of equity returns, the proposed method achieves a more favorable bias–variance trade-off than purely analytical cleaners and delivers systematically lower out-of-sample cross–covariance prediction errors. Our results demonstrate that combining random-matrix theory with machine learning makes asymptotic theories practically effective in realistic time-varying markets.
Keywords Spectral Tokenization $\cdot$ Random Matrix Theory $\cdot$ Equivariant Neural Networks $\cdot$ Nonlinear Shrinkage $\cdot$ Non-Stationary Dependence
1 Introduction
Empirical cross–covariance matrices provide a compact way to summarize how two groups of variables move together. Consider two zero–mean random vectors, $\mathbf{X}\in\mathbb{R}^{n_{x}}$ and $\mathbf{Y}\in\mathbb{R}^{n_{y}}$, observed over a time window of length $\Delta t_{\textrm{in}}$. Their cross–covariance matrix $\mathbf{\Sigma}_{XY}$ contains the expected products between components of $\mathbf{X}$ and $\mathbf{Y}$, and can therefore be interpreted as a linear measure of co-movement. It indicates which fluctuations in one set of variables tend to be associated with fluctuations in the other. In empirical applications, $\mathbf{\Sigma}_{XY}$ is not known and must be estimated from the $\Delta t_{\textrm{in}}$ paired observations available in the window.
A fundamental challenge arises when the number of variables is not small compared to the window length. In such high–dimensional regimes, the empirical estimator is heavily affected by sampling noise where apparent relationships may emerge purely by chance, and the resulting estimate can behave unreliably when evaluated Out-Of-Sample (OOS) [9]. For standard covariance matrices, this issue has been studied extensively, and a large body of work has developed principled denoising strategies, often based on spectral methods and Random Matrix Theory (RMT), to separate stable structure from noise [7]. The cross–covariance setting is less straightforward. When $n_{x}\neq n_{y}$, the matrix $\mathbf{\Sigma}_{XY}\in\mathbb{R}^{n_{x}\times n_{y}}$ is rectangular, so the usual eigenvalue-based tools do not apply directly. Instead, one must work with a singular value decomposition. While singular-value filtering provides a natural analogue of covariance cleaning, the corresponding theoretical framework for choosing optimal shrinkage rules is comparatively less mature, and its effectiveness, in realistic time-varying systems, remains an open question.
Related denoising results exist for additive-noise models [17, 11], but sample (cross-)covariances arise from products of observations and therefore call for a multiplicative-noise cleaning framework. To the best of our knowledge, the only available analytical characterization of an asymptotically optimal singular-value shrinkage for empirical cross-covariance matrices under multiplicative noise is due to Benaych-Georges, Bouchaud, and Potters (BBP) [2]. In the high-dimensional asymptotic regime, they derive a Rotationally Invariant Estimator (RIE) that preserves the empirical singular vectors of $\widehat{\mathbf{\Sigma}}_{XY}$ while shrinking its singular values so as to minimize the asymptotic Mean Squared Error (MSE) relative to the true cross-covariance $\mathbf{\Sigma}_{XY}$. Their result is presented in two versions: a baseline setting with identity marginals ($\mathbf{\Sigma}_{XX}=\mathbf{I}_{n_{x}}$ and $\mathbf{\Sigma}_{YY}=\mathbf{I}_{n_{y}}$), which also covers the practically relevant case of whitened variables, and a more general setting that incorporates non-trivial marginal covariances into the optimal correction. The theory is validated through simulations and controlled benchmarks, but it does not consider real financial data and relies on strong stationarity assumptions that are rarely met in practice.
The impact of non-stationarity is already well documented in the simpler covariance case. Bongiorno, Challet and Loper [3] show that covariance–shrinkage estimators derived under stationarity assumptions can underperform naive asset–agnostic shrinkages when tested on real markets. In particular, they report that shrinkage-based portfolios may be dominated OOS by simple strategies such as the Average Oracle, which is calibrated on outdated returns and yet evaluated on more recent data, and that this behaviour is robust across different datasets and markets.
In the covariance setting, the mismatch between analytically optimal shrinkage rules and the empirical non-stationarity of financial data has been mitigated through physics-informed neural network architectures that embed the random-matrix solution as a limiting case [5, 6]. In these approaches, the RMT shrinkage is recast as a constrained network that remains expressive enough to recover the asymptotically optimal filter under stationarity, while gaining flexibility via additional trainable degrees of freedom to operate beyond that regime. A key element is the explicit enforcement of the problem’s symmetries, notably the rotationally invariant structure implied by RMT. Trained on rolling windows, the model learns to interpolate between a denoising behavior effective in quasi-stationary periods and a predictive correction that captures persistent time variation, yielding improved OOS performance relative to purely analytical shrinkage and naive baselines. Moreover, the neural formulation naturally accommodates decision-aware training. Rather than optimizing an MSE on covariance entries, one can directly minimize an objective aligned with the downstream portfolio loss [4, 13], thereby linking statistical estimation to the relevant economic criterion.
In this work, we introduce a physics-informed learning framework for cross–covariance estimation that preserves the rotational invariances of RMT while operating beyond the stationarity and boundedness regimes required by existing analytical shrinkage. The key idea is to parameterize the cleaned estimator directly in the empirical singular-vector basis and to learn a dimension-agnostic nonlinear map from empirical singular values and marginal projections to cleaned singular values, so that the BBP correction is recovered as a limiting case, but additional degrees of freedom can accommodate macroscopic modes and time variation. Empirically, we observe that the analytically optimal BBP shrinkage can lose robustness on equity returns as the universe size increases in the presence of a dominant market mode, whereas the proposed symmetry-preserving estimator remains stable and yields lower OOS reconstruction error, including in OOS matrix sizes.
2 Two-Stream Architecture for Cross–Covariance Cleaning
We start from the empirical cross-correlation $\widehat{\mathbf{C}}_{XY}\in\mathbb{R}^{n_{x}\times n_{y}}$ computed on a window of length $\Delta t_{\textrm{in}}$, together with the marginal empirical correlations $\widehat{\mathbf{C}}_{XX}$ and $\widehat{\mathbf{C}}_{YY}$. Let $r:=\min(n_{x},n_{y})$ and write the Singular Value Decomposition (SVD)
$\widehat{\mathbf{C}}_{XY}=\sum_{k=1}^{r}\widehat{s}_{k}\,\widehat{\mathbf{u}}_{k}\widehat{\mathbf{v}}_{k}^{\top}.$ (1)
When needed, we complete $\{\widehat{\mathbf{u}}_{k}\}_{k=1}^{r}$ and $\{\widehat{\mathbf{v}}_{k}\}_{k=1}^{r}$ to orthonormal bases of $\mathbb{R}^{n_{x}}$ and $\mathbb{R}^{n_{y}}$; the extra directions correspond to zero singular values. A RIE keeps the empirical singular vectors $\{\widehat{\mathbf{u}}_{k},\widehat{\mathbf{v}}_{k}\}$ and replaces each $\widehat{s}_{k}\mapsto\widetilde{s}_{k}$ by a cleaned value.
Motivated by the structure of BBP shrinkage [2], the correction should depend on the empirical singular spectrum and on how marginal energy projects onto the corresponding singular directions. We therefore define the marginal projections
$\widehat{\gamma}^{(x)}_{k}:=\widehat{\mathbf{u}}_{k}^{\top}\widehat{\mathbf{C}}_{XX}\widehat{\mathbf{u}}_{k},\qquad k=1,\ldots,n_{x},\qquad\qquad\widehat{\gamma}^{(y)}_{k}:=\widehat{\mathbf{v}}_{k}^{\top}\widehat{\mathbf{C}}_{YY}\widehat{\mathbf{v}}_{k},\qquad k=1,\ldots,n_{y},$ (2)
and the aspect ratios $q_{x}:=n_{x}/\Delta t_{\textrm{in}}$ and $q_{y}:=n_{y}/\Delta t_{\textrm{in}}$. When $n_{x}\neq n_{y}$, only the first $r$ singular values are nonzero, while the larger side retains additional marginal directions that enter BBP-type functionals.
To expose this information as a single aligned sequence, we set $p:=\max(n_{x},n_{y})$ and define padded scalars for $k=1,\ldots,p$,
$\overline{s}_{k}:=\begin{cases}\widehat{s}_{k},&k\leq r,\\ 0,&k>r,\end{cases}\qquad\overline{\gamma}^{(x)}_{k}:=\begin{cases}\widehat{\gamma}^{(x)}_{k},&k\leq n_{x},\\ 0,&k>n_{x},\end{cases}\qquad\overline{\gamma}^{(y)}_{k}:=\begin{cases}\widehat{\gamma}^{(y)}_{k},&k\leq n_{y},\\ 0,&k>n_{y}.\end{cases}$ (3)
[…]
For each $k\in\{1,\ldots,p\}$ we build two stream-specific tokens
$\boldsymbol{\tau}^{(x)}_{k}:=\bigl[\overline{\gamma}^{(x)}_{k},\ \overline{s}_{k},\ q_{x}\bigr]^{\top},\qquad\boldsymbol{\tau}^{(y)}_{k}:=\bigl[\overline{\gamma}^{(y)}_{k},\ \overline{s}_{k},\ q_{y}\bigr]^{\top}.$ (4)
A shared encoder $E_{\theta}$, implemented as a token-wise two-layer Multi-Layer-Perceptron (MLP) with a 16-unit hidden layer and LeakyReLU activation, maps each token to a 2-dimensional embedding
$\mathbf{e}^{(x)}_{k}=E_{\theta}(\boldsymbol{\tau}^{(x)}_{k})\in\mathbb{R}^{2},\qquad\mathbf{e}^{(y)}_{k}=E_{\theta}(\boldsymbol{\tau}^{(y)}_{k})\in\mathbb{R}^{2}.$ (5)
Then, we fuse the two streams additively,
$\mathbf{z}_{k}=\mathbf{e}^{(x)}_{k}+\mathbf{e}^{(y)}_{k},\qquad k=1,\ldots,p,$ (6)
and pass the sequence $(\mathbf{z}_{1},\ldots,\mathbf{z}_{p})$ through a two-layer bidirectional LSTM aggregator (hidden sizes 128 and 64). This yields context-aware states $(\mathbf{h}_{1},\ldots,\mathbf{h}_{p})$ that provide each index with global spectral context [5].
A pointwise head $g_{\theta}$, implemented as a token-wise two-layer MLP with a 252-unit hidden layer and LeakyReLU activation, maps each state to a scalar residual correction $\delta_{k}=g_{\theta}(\mathbf{h}_{k})$. We keep only the first $r$ corrections and define the cleaned singular values by the additive parametrization
$\widetilde{s}_{k}=\widehat{s}_{k}+\delta_{k},\qquad k=1,\ldots,r.$ (7)
See Appendix 5 for a step-by-step derivation showing that BBP is recovered as a special case.
Alternatively, one may enforce a bounded multiplicative correction via $\widetilde{s}_{k}=\widehat{s}_{k}\,\sigma(\delta_{k})$, where $\sigma(\cdot)$ denotes the logistic sigmoid, which guarantees $0\leq\widetilde{s}_{k}\leq\widehat{s}_{k}$ (see Appendix 5.3 and 8.4 for more details). In our experiments, however, this constrained parametrization led to slightly more difficult optimization, and we therefore adopt the additive form. To provide a clearer conceptual overview of the integrated components and the associated data flow, we illustrate the architecture in Figure 1.
Figure 1: Flowchart of the neural singular value cleaning architecture. The diagram illustrates the construction of dual-stream tokens from marginal projections $\overline{\gamma}$ and singular values $\overline{s}$, their transformation through a shared encoder $E_{\theta}$, and the global context aggregation via a bidirectional LSTM and pointwise head $g_{\theta}$ to produce the final additive corrections $\delta_{k}$.
Finally, we reconstruct the cleaned cross-covariance by plugging Eq. (7) in Eq. (1) and rescaling with the marginal sample standard deviations,
$\widetilde{\boldsymbol{\Sigma}}_{XY}=\widehat{\boldsymbol{D}}_{X}\widetilde{\mathbf{C}}_{XY}\widehat{\boldsymbol{D}}_{Y}.$ (8)
In this work, we train in the cross-correlation domain to isolate multivariate filtering, but the same pipeline can be optimized end-to-end for cross-covariances by including the rescaling in the objective. For supervision, given an OOS target $\mathbf{C}^{\mathrm{OOS}}_{XY}$ computed on the subsequent window of length $\Delta t_{\mathrm{out}}$, we minimize the MSE, i,e., the squared Frobenius loss
$\mathcal{L}(\theta)=\frac{1}{n_{x}n_{y}}\,\bigl\|\widetilde{\mathbf{C}}_{XY}-\mathbf{C}^{\mathrm{OOS}}_{XY}\bigr\|_{F}^{2}.$ (9)
The network has 331,355331{,}355 learnable parameters, independent of $(n_{x},n_{y},\Delta t_{\textrm{in}})$. It can be trained at one range of dimensions and deployed at other sizes without retraining. Moreover, it is equivariant to permutations of the $X$ and $Y$ coordinates. A Python implementation is available on GitHub [10]
3 Validation on Synthetic and Market Data
To assess our Neural Network (NN) based estimator, we first benchmark it on synthetic datasets that reproduce the spectral regimes of finite-rank signals and heavy-tailed noise under which the BBP shrinkage is derived [2]. These experiments therefore constitute a favorable setting for the analytical method, relying on stylized assumptions such as identity auto-covariance blocks that are mathematically convenient yet seldom satisfied in financial data. As reported in Table1, our estimator matches the BBP performance across these validation tests, indicating that, in stationary regimes where the theoretical assumptions hold, the network can recover the optimal rotationally invariant filtration without introducing spurious structure. We further add a fourth benchmark in which a macroscopic one-factor mode is injected into the population correlation matrix $\boldsymbol{C}$, thereby violating the BBP boundedness assumption [2]; in this regime, BBP performance degrades. To confirm that this deterioration is indeed driven by the lack of boundedness, we also designed a Monte Carlo Cross-Validated (CV) BBP variant, which remains robust under the same perturbation. For details on the boundedness issue, the CV procedure, and the data-generating processes and validation protocol, we refer the reader to Appendix 6 and 7.
Condition Empirical (MLE) Analytical (BBP) Numerical (CV) Neural Network (NN)
Panel A: Finite-Rank Spiked Models (varying spike % $\xi$)
$\xi=40\%$ $2.00\times 10^{-3}$ $2.46\times 10^{-4}$ $\mathbf{2.46\times 10^{-4}}$ $2.53\times 10^{-4}$
$\xi=30\%$ $2.00\times 10^{-3}$ $1.92\times 10^{-4}$ $\mathbf{1.92\times 10^{-4}}$ $1.95\times 10^{-4}$
$\xi=20\%$ $2.00\times 10^{-3}$ $\mathbf{1.34\times 10^{-4}}$ $\mathbf{1.33\times 10^{-4}}$ $1.34\times 10^{-4}$
$\xi=10\%$ $2.00\times 10^{-3}$ $\mathbf{7.03\times 10^{-5}}$ $\mathbf{7.00\times 10^{-5}}$ $7.12\times 10^{-5}$
$\xi=0\%$ $2.00\times 10^{-3}$ $\mathbf{3.73\times 10^{-6}}$ $3.92\times 10^{-6}$ $9.02\times 10^{-6}$
Panel B: Heavy-Tailed Bulk Models (varying tail exponent $\alpha$)
[…]
$\alpha=1.5$ $2.00\times 10^{-3}$ $\mathbf{6.19\times 10^{-4}}$ $6.20\times 10^{-4}$ $\mathbf{6.18\times 10^{-4}}$
Panel C: White Heavy-Tailed Models (varying tail exponent $\alpha$)
Gaussian $2.00\times 10^{-3}$ $\mathbf{5.63\times 10^{-4}}$ $5.64\times 10^{-4}$ $5.64\times 10^{-4}$
$\alpha=5.0$ $2.00\times 10^{-3}$ $\mathbf{5.63\times 10^{-4}}$ $5.65\times 10^{-4}$ $5.64\times 10^{-4}$
$\alpha=2.5$ $2.00\times 10^{-3}$ $\mathbf{2.50\times 10^{-4}}$ $\mathbf{2.50\times 10^{-4}}$ $\mathbf{2.51\times 10^{-4}}$
$\alpha=1.5$ $2.00\times 10^{-3}$ $\mathbf{3.92\times 10^{-5}}$ $\mathbf{3.90\times 10^{-5}}$ $4.14\times 10^{-5}$
Panel D: Gaussian Bulk Models with Common Mode (varying mode $m$)
$m=0.5$ $1.09\times 10^{-3}$ $7.90\times 10^{-2}$ $\mathbf{7.45\times 10^{-4}}$ $\mathbf{7.42\times 10^{-4}}$
$m=0.4$ $1.42\times 10^{-3}$ $3.29\times 10^{-2}$ $\mathbf{9.25\times 10^{-4}}$ $\mathbf{9.18\times 10^{-4}}$
$m=0.3$ $1.66\times 10^{-3}$ $9.89\times 10^{-3}$ $\mathbf{9.83\times 10^{-4}}$ $\mathbf{9.84\times 10^{-4}}$
$m=0.2$ $1.85\times 10^{-3}$ $2.23\times 10^{-3}$ $\mathbf{9.84\times 10^{-4}}$ $\mathbf{9.79\times 10^{-4}}$
$m=0.1$ $1.96\times 10^{-3}$ $8.85\times 10^{-4}$ $\mathbf{8.52\times 10^{-4}}$ $\mathbf{8.56\times 10^{-4}}$
Table 1: Summary of the average MSE for different estimators across three synthetic regimes over 1000 simulations. Bold marks the lowest MSE; ties are bold when a 95% percentile $100,000$ copies bootstrap cannot distinguish them. Panel A varies the spike percentage $\xi$ (Finite-Rank). Panel B and C vary the tail exponent $\alpha$ (Heavy-Tailed Bulk and White models), where a lower $\alpha$ implies heavier tails. Panel D is a variation of Panel B (Gaussian) with a one-factor mode injected. All simulations are performed with $n_{x}=200$, $n_{y}=350$, $\Delta t_{\textrm{in}}=500$ [2].
Moving to real-world data, we train and test the estimators on a universe of the top 1000 U.S. equities by market capitalization (1995-2024), filtered for liquidity and stability (see Appendix 8.1). To mitigate look-ahead bias, the top-1000 universe is reconstituted at each window start, so constituents vary over time. In this regime, the non-stationarity of financial time series becomes a predominant factor. According to standard practice, we train a separate model for each calendar year and then test it on the subsequent two years, using as a target a realized OOS correlation matrix computed over a hold-out window of length $\Delta t_{\mathrm{out}}=240$ trading days. Each yearly model is optimized for 50 epochs (500 gradient steps per epoch) with mini-batches of 32 samples. Each sample is generated by drawing a random date, the total number of assets $n:=n_{x}+n_{y}\sim\mathcal{U}[50,500]$, and the relative dimension $\nu:=n_{x}/(n_{x}+n_{y})\sim\mathcal{U}[0.05,0.95]$, which enforces dimension-agnostic behavior and allows evaluation on larger shapes ($n>500$) without retraining.
Figure 2: OOS MSE for cross-correlation estimators on financial data covering 2017–2024. For each test window, the estimator is trained on an expanding sample from 1995 up to the year immediately preceding the test period, and the reported value averages 1,000 independent runs for each out-of-sample year. The figures report also the Oracle (red dashed line), which is the lower achievable MSE by a RIE (see eq. (35)). The left panels preserve the chronological order between in-sample and out-of-sample windows, whereas the right panels shuffle dates. The upper panels varies the total number of assests for $\nu=0.25$, the lower panels varies the relative dimension $\nu$ for $n=1000$. The error bars indicate a $95\%$ percentile $100,000$ copies bootstrap confidence interval.
Figure 2 (upper left) reports the OOS MSE as a function of the total number of assets $n$ for a fixed in-sample window ($\Delta t_{\textrm{in}}=500$). The Maximum-Likelihood Estimator (MLE, yellow) provides the unfiltered baseline. Across both the $n$- and $\nu$-sweeps, the proposed NN estimator (blue) achieves the lowest error. For small universes (e.g., $n=250$), its performance is not statistically distinguishable from BBP (gray). However, as $n$ increases, BBP deteriorates sharply and can even underperform MLE, indicating that the analytical shrinkage can become harmful in this regime. We attribute this failure to the presence of a macroscopic market mode in real returns, which violates the boundedness conditions underlying BBP. This interpretation is supported by two additional diagnostics. First, the numerical CV estimator is not affected by the same degradation and remains comparable to the NN. Second, in Appendix 8.3 we show that pre-processing the data by removing the mode, operationally defined as the daily cross-sectional mean return, eliminates the gross of the BBP underperformance. Finally, when we shuffle the time series to break potential non-stationarity between in-sample and out-of-sample periods, BBP remains unstable whereas CV continues to track the NN, further indicating that the dominant driver is the mode rather than temporal non-stationarity, see Fig. 2 (right) . Notably, the NN performance remains robust even for $n>500$, which constitutes a fully out-of-distribution regime for the model. This region lies beyond the matrix sizes observed during training ($n\leq 500$), and it also corresponds to a different historical period than the data used to fit the network. To further probe this extrapolative regime, in the lower panel of Fig. 2 (lower) we unpack the $n=1000$ case by varying the relative dimension $\nu$, and we find that the NN gain remains stable across the full range of $\nu$, indicating that its advantage is not concentrated in a specific configuration.
Our working hypothesis is that this behavior is precisely the one targeted by the proposed architecture. In stationary settings, the network should primarily act as a noise-cleaning operator on the empirical spectrum. In contrast, when applied to non-stationary financial data, it can interpolate between denoising and a learned predictive correction that exploits persistent structure beyond sampling noise. To confirm that this gain is derived from capturing genuine temporal signals (non-stationarity), we replicate the analysis on temporally shuffled data. Shuffling destroys the serial correlations responsible for the predictive edge, creating a purely stationary environment. Figures 2 (right) illustrate that performance degradation of BBP persists even in the shuffled regime, confirming the underlying structural driver. As expected, the CV estimator does not degrade and instead matches the NN performance indicating that in a stationary regime the NN learn how to denoise only sample noise. This interpretation is further supported by the feasibility diagnostic reported in Appendix 8.4.
4 Discussion and Conclusion
Across controlled synthetic benchmarks and temporally shuffled equity returns, the proposed estimator matches the performance of the optimal rotationally invariant shrinkage, indicating that it has learned a stationary random-matrix-consistent spectral denoiser rather than introducing spurious structure. In the presence of a pronounced market mode, it also improves upon the corresponding analytic shrinkage method, which is consistent with a mode-aware generalization.
On original (unshuffled) market data, the two cleaners remain comparable for moderate universes, but their behavior diverges as $n$ grows. The analytical BBP shrinkage deteriorates sharply with system size and can even underperform the unfiltered MLE, consistent with sensitivity to macroscopic components (notably a market mode) that violate the boundedness conditions underlying the asymptotics. In contrast, the neural estimator remains stable and achieves lower OOS reconstruction error, including for $n>500$ (an out-of-distribution regime relative to training). Removing the market mode substantially reduces the BBP collapse, yet a residual gap to the neural estimator persists in the chronological evaluation and largely disappears under temporal shuffling; taken together, these controls support the view that the architecture preserves the BBP symmetry constraints while using OOS supervision to retain only cross-dependencies that persist across adjacent windows, i.e., a symmetry-preserving singular-value map that interpolates between denoising and a mild predictive correction.
A natural direction for future work is to extend the same symmetry-preserving parametrization to lead-lag cross-covariance matrices, where the relevant signal is intrinsically dynamical and cannot be reduced to a single synchronous estimator. In this setting, rather than training via the Frobenius loss, one can adopt an end-to-end decision-aware objective consistent with a trading strategy so that the shrinkage rule is learned under the same decision loss used for evaluation.
Acknowledgments
This work was performed using HPC resources from the “Mésocentre” computing center of CentraleSupélec and École Normale Supérieure Paris-Saclay supported by CNRS and Région Île-de-France (mesocentre.centralesupelec.fr). E.M. acknowledges a fellowship funded by PNRR for the PhD DOT1608375 in Sistemi complessi per le scienze fisiche, socio-economiche e della vita of Catania University.